Randomly selected birth records were​ obtained, and categorized as listed in the table to the right. Use a 0.01 significance level to test the reasonable claim that births occur with equal frequency on the different days of the week. How might the apparent lower frequencies on Saturday and Sunday be​ explained?

DAY SUN MON TUES WED THURS FRI SAT
NUMBER OF BIRTHS 39 56 62 55 55 61 50

1. Determine the null and alternative hypotheses.

Ho =

H1 =

2. Calculate the test​ statistic x^2 = ___________________ ​(Round to three decimal places as​ needed.)

3. Calculate the​ P-value. ​P-value = __________________ ​(Round to four decimal places as​ needed.)

4. What is the conclusion for this hypothesis​ test?

A. Reject Upper H 0. There is sufficient evidence to warrant rejection of the claim that births occur with equal frequency on the different days of the week.

B. Fail to reject Upper H 0. There is sufficient evidence to warrant rejection of the claim that births occur with equal frequency on the different days of the week.

C. Fail to reject Upper H 0. There is insufficient evidence to warrant rejection of the claim that births occur with equal frequency on the different days of the week.

D. Reject Upper H 0. There is insufficient evidence to warrant rejection of the claim that births occur with equal frequency on the different days of the week.

5. How might the apparent lower frequencies on Saturday and Sunday be​ explained?

A. Induced or​ Caesarean-section births are scheduled on weekends whenever possible.

B. Induced or​ Caesarean-section births are scheduled during the week whenever possible.
Answer #1

1. Ho = nullhypothesis = The number of births on any day is equally likely i.e. b(mon) = b(tue) = b(wed) = b(thu) = b(fri) = b(sat) = b(sun)

Ha = alternate hypothesis = The number of births on any day is not equally likely

2. chi square test statistic = 6.667

3. Corresponding p value for 6 degrees of freedom = 0.353

4.

C. Fail to reject Upper H 0. There is insufficient evidence to warrant rejection of the claim that births occur with equal frequency on the different days of the week.

5.

B. Induced or​ Caesarean-section births are scheduled during the week whenever possible.